Explain why the sin(theta) and cos(theta) are periodic functions?
1
Expert's answer
2011-02-21T06:26:31-0500
From the definition of the periodic function,
f(x) = f(x+kT), where k=0,1,2,3,... and T is the period.
cos (θ) = cos(θ +k 2π)
sin(θ) = sin(θ + k 2π)
2π is the period of cos and sin.